OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 – 5x2 + 6x ?

OpenStudy (loser66):

factor x out first, tell me what you have

OpenStudy (anonymous):

what do you mean ?

OpenStudy (loser66):

oh, sorry, I don't know what I mean, too. let me tag other for help. @Jhannybean

OpenStudy (jhannybean):

Fator out an x first. you see that there is a common "x" multiplier in this function, so pull out out of all the terms. You would have \[\large f(x) = x^3 - 5x^2 +6x \]\[\large f(x) = x(x^2 -5x +6)=0\] Can you finish this up now? :)

OpenStudy (anonymous):

What do i do ? :()

OpenStudy (jhannybean):

what are two numbers that multiply to get 6, but add to give you -5?

OpenStudy (jhannybean):

hint: think of the multiples of 6.

OpenStudy (anonymous):

(x-3)(x-2)

OpenStudy (jhannybean):

good. so now you have 3 terms to solve for. x = 0 x - 3 = 0 x - 2 = 0 Can you find the zeroes of your function?

OpenStudy (anonymous):

so 3 and 2 ?

OpenStudy (jhannybean):

and 1 more

OpenStudy (anonymous):

oopss 0 , 3 , 2

OpenStudy (jhannybean):

Yay! you got it.

OpenStudy (anonymous):

ok . one more ?

OpenStudy (anonymous):

According to The Fundamental Theorem of Algebra, how many zeros does the function f(x) = 24x5 + 8x3 – 2x – 15 have?

OpenStudy (jhannybean):

To find this, analyze the variable with the highest power.

OpenStudy (anonymous):

24

OpenStudy (jhannybean):

no the power, not the coefficient. \[\large f(x) =24\color{green}{x^5}+8x^3-2x-15 \]

OpenStudy (anonymous):

so its 5 ?

OpenStudy (anonymous):

hello ?

OpenStudy (jhannybean):

Yes, 5. ) Good job.

OpenStudy (jhannybean):

:)*

OpenStudy (anonymous):

tbhankyoouu (:

OpenStudy (jhannybean):

npppp :)

OpenStudy (jhannybean):

and THANK YOU for the medal!